OpenStudy (anonymous):

V varies inversely with T and V =18 when T= 3. Which equation shows this relationship?

Nnesha (nnesha):

\[\huge\rm y = \frac{ k }{ x }\] formula for direct variation where y represent V and x represent T replace t and v by their values

OpenStudy (ribhu):

v = k/T

Nnesha (nnesha):

inverse variation**

Nnesha (nnesha):

there is a mistake inverse** not direct

OpenStudy (ribhu):

v = 54/T

Nnesha (nnesha):

at least give him/her chance to solve it